Evaluate f(x) = 3x5 – 5x3 + 2x2 + 6 for:
a. x = 0
b. x = 2
c. x = -2
d. 2x

Answer by scianci(186)

a. f(0) = 6 [all the terms containing an x are canceled out]
b. f(2) = 3(2)^5 - 5(2)^3 + 2(2)^2 + 6
= 3(32) - 5(8) + 2(4) + 6
= 96 - 40 + 8 + 6
= 70
c. f(-2) = 3(-2)^5 - 5(-2)^3 + 2(-2)^2 + 6
= 3(-32) - 5(-8) + 2(4) + 6
= -96 + 40 + 8 + 6
= -42
d. f(2x) = 3(2x)^5 - 5(2x)^3 + 2(2x)^2 + 6
= 3(32x^5) - 5(8x^3) + 2(4x^2) + 6
= 96x^5 - 40x^3 + 8x^2 + 6
[it is not possible to give a numerical answer for this one unless you have a specific value of x to plug in]